A 300kg block start at rest from the top of 30° incline and accelerate uniformly down the incline ,moving 20m in 1.50s. (a)Find the magnitude of the acceleration of the block .(b)find the coefficient of friction between the block and the incline .(c)Find the magnitude of the frictional force acting on the block.(d)Find the speed of the block after it has slide a distance of2.00m.[g=9.8ms squared]




1
Expert's answer
2021-12-06T09:46:46-0500

*We assume here that the distance is 2 m, not 20 m.

(a) The acceleration is


"a=\frac{2d}{t^2}=1.78\text{ m/s}^2."


(b) The coefficient of friction:


"\mu=\tan\theta-\frac{a}{g\cos\theta}=0.368."


(c) The frictional force:


"f=\mu N=\mu mg\cos\theta=937\text{ N}."

(d) The speed is:


"v=\sqrt{2ad}=2.67\text{ m/s}."